The enthalpy and entropy change of a reaction are -3.9 kJ/mole and +56.6

J/mole K respectively at 25o C. What is the free energy change in kJ/mole?

Is this reaction always spontaneous, never spontaneous or does it depend on

the temperature?


1
Expert's answer
2021-11-18T06:06:02-0500

∆G = ∆H - ∆TS

∆H = -3.9KJ/mol

T = 25 + 273 = 298K

∆S = 56.6J/mol = 0.056KJ/mol

∆G = -3.9 - 298(0.056)

= -3.9 - 16.688

= -20.588KJ/mol

The reaction is spontaneous